# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v38

@@ 1 Introduction
Despite this robust evidence of joint price formation across option contracts, existing informed-trading models (e.g., Back (1993b); Biais and Hillion (1994); Brennan and Cao (1996); Easley et al. (1998b); Collin-Dufresne et al. (2021)) remain confined to single-option settings and therefore cannot speak to the cross-contract mechanisms at the core of these empirical findings—cross-strike dynamics, implied-volatility-surface formation, and multi-option trades such as the straddle. Our model yields cross-option equilibrium restrictions that sharpen and generalize the empirical intuition while generating additional testable implications. For instance, we identify cross-strike volume imbalance as the equilibrium channel behind the return predictability of signed option volume documented by Pan and Poteshman (2006), and provide the first equilibrium rationale for the volatility-forecasting power of straddle demand observed in Ni et al. (2008) and Goyal and Saretto (2009).
@@
Empirical work consistently finds that access to derivatives markets reduces information asymmetry and improves price efficiency—evidence spans futures (e.g., Garbade and Silber (1983); Chan (1992)), credit derivatives (e.g., Blanco et al. (2005); Longstaff et al. (2005); Acharya and Johnson (2007)), as well as options (e.g., Hu (2018)). Theoretical foundations for this efficiency channel, however, remain limited. Our model isolates a simple and intuitive mechanism: once derivatives are tradable, an informed trader is no longer restricted to (generally suboptimal) linear demand and can instead express his information through targeted state-contingent positions. In other words, we show that the linear equilibrium in Kyle (1985) is not robust to the introduction of derivatives. [^4] See Example 7.1. Derivatives thus expand the information that prices can aggregate beyond what the underlying asset alone can support.
@@
The reduction of Section 6 is the key modeling innovation that delivers tractability at our full Arrow-Debreu generality and yields an equilibrium characterized by a single endogenous constant. Readers who prefer the high-level overview without the formal developments can see the reduction distilled in Example 6.7, which links directly back to the simplified environment of Section 2.
@@ 2 Basic Intuition
In choosing the scale $\alpha$ of his long–short portfolio, the insider trades off the marginal payoff from exploiting the true payoff distribution against the marginal cost of trading. That cost has two components. First, there is the cost of executing at prevailing prices, absent any price impact. Second, there is an informational component: larger trades make order flow more informative for the market maker, which amplifies cross-market price impact and erodes the insider’s informational advantage. The optimal scale $\alpha^{*}$ balances these margins (see Figure 1a). Equilibrium is characterized by this single endogenous constant—nontrivially, this same equilibrium structure carries over to the general model (see Example 6.7). Allowing correlated noise trades, varying noise intensity, and non-uniform priors has no material effect on the main results. [^7] Correlated Noise Trades Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has an opposite sign to their payoff correlation. In that case, the insider would scale up his trading strategy to exploit the noise traders’ additional losses from trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin remains the same—akin to having biased noise trades in the single-asset setting.  [^8] Varying Noise Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}.$ Thus, as expected, higher noise trading intensity reduces price impact both within each and across markets.  [^9] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{{\rm e}^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{{\rm e}^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$ by causing the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
@@
If fewer than all three AD securities are available for trade, both the insider’s profit and the information efficiency of prices decline. Here, information efficiency is naturally measured by the posterior probability weight that the market maker assigns to the true signal. Derivative markets allow informed traders’ orders to better express (and therefore reveal) their information, thereby its incorporation into prices. Therefore, derivatives do more than “add leverage”: they change what information can be traded and how that information propagates across markets. This result provides an equilibrium foundation for a long-held intuition that derivatives concentrate informed trading in directions that are difficult to trade in the underlying alone (see Black (1975)), and it is consistent with evidence that option markets contain incremental information about underlying returns and higher moments (see Easley et al. (1998a); Pan and Poteshman (2006); Gârleanu et al. (2009)).
Example: “Straddle”
A key implication is not merely that derivative trades reflect information, but that that information is transmitted through cross-market price impact. We demonstrate this mechanism in the case of volatility. Interpret the three states $x_{1}<x_{2}<x_{3}$ as the possible $t=1$ realizations of an underlying asset price, and normalize the $t=0$ price to be $x_{2}$. Then $s_{1}$ is the “high volatility” signal and $s_{2}$ the “low volatility” signal. The signal $s_{1}$ reveals that the asset price will move, but without a specified direction, whereas $s_{2}$ reveals that there will be no price movement. Security $x_{1}$ is a put-like claim (it pays when the asset price falls), while $x_{3}$ is call-like (it pays when the asset price rises). If the insider observes $s_{1}$, he buys the put-call pair $(x_{1},x_{3})$—a “straddle”—to capitalize price movement regardless of direction. In real option markets, the straddle is a standard volatility strategy (see Hull (2015)) and is widely discussed in the empirical literature. [^10] See Coval and Shumway (2001); Pan and Poteshman (2006); Ni et al. (2008); Driessen et al. (2009); Goyal and Saretto (2009). This strategy, along with other common option strategies, will be derived as an equilibrium portfolio in the general model.
@@ 3 General Model
We place no restrictions on $\eta(\,\cdot\,|\,\cdot\,)$, i.e., no restrictions on the probability distributions across states or the insider’s information.
@@
A continuous normal prior and payoff distribution, as in Kyle (1985), can be approximated arbitrarily well by discrete state-signal grids. Accordingly, our results do not depend on whether the state and signal spaces are discrete or continuous. We work with the discrete formulation here to minimize technicality. A fully rigorous continuous-space treatment (with general signal structures and priors) is provided in Keller and Tseng (2026).
In this single-asset special case, the price impact is the slope $\lambda$ of the market maker’s inverse supply function $P(Q)=v_{0}+\lambda Q$, as specified by Equations (5) and (6). Price impact is proportional to the noise-adjusted signal variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$. Higher signal variation leads to greater informed demand variation across signals, making order flow more informative for the market maker and resulting in a higher price impact and reduced informed demand. Thus, the price impact and informed demand are determined by the information intensity of the insider’s signal. As suggested in Section 2, this intuition generalizes across AD markets.
@@
Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. In this case, any portfolio $W(\,\cdot\,)$ of AD securities can be replicated using a combination of the underlying asset and options: [^13] See Back (2010), Exercise 3.5. Let $x_{i_{0}}$ be the expected $t=1$ asset price under the market maker’s prior at $t=0$. For any $W(\,\cdot\,)$, we can find $(a_{i})$ such that
$\!\!\!W(x_{j})$ $=a_{i_{0}}(x_{j}-x_{i_{0}})+\sum_{i<i_{0}}a_{i}(x_{j}-x_{i})_{-}+\sum_{i>i_{0}}a_{i}(x_{j}-x_{i})_{+},\quad\mbox{for $j=1,2,\cdots$}.$ (7)
That is, $W(\,\cdot\,)$ can be replicated by holding $a_{i_{0}}$ shares in the underlying asset, and $a_{i}$ shares in out-of-the-money put (resp. call) options at strikes $x_{i}<x_{i_{0}}$ (resp. $x_{i}>x_{i_{0}}$). Thus, the AD markets can be replaced by an equivalent set of asset and option markets.
The insider learns the true payoff distribution $\eta(\,\cdot\,|s_{k})$ of the underlying asset from his private signal $s_{k}$ and then submits orders for the underlying asset and options. Noise trades follow $\varepsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ in each market $i$ (for the asset or option at strike $x_{i}$). When the insider submits orders $(a_{i})$, the order flow received by the market maker is given by $a_{i}+\varepsilon_{i}$, for $i=1,2,\cdots.$
The rest of the model proceeds along the same lines as the AD formulation. Conversely, the results obtained in the AD formulation translate to options.
@@
The options formulation relates to the single-asset case as follows:
@@
(i) If the payoff distributions differ across signals only in their expected values, and there are no option markets, then our model reduces to the Kyle (1985) framework, where only the underlying asset is traded.
(ii) Conversely, if all possible payoff distributions have the same expected value, then the insider has no incentive to trade the underlying asset. In other words, exploiting information beyond the expected payoff requires options.
Proposition 3.3(i) applies only when there are no option markets. In other words, the linear equilibrium (5) in Kyle (1985) is not robust to the introduction of options. Options allow the insider to optimally scale back his demands for low-probability strikes, resulting in an overall demand that is inherently nonlinear. This point will be illustrated explicitly in Example 7.1 below.
@@ 4 Cross-Market Inference
Intuitively, the market maker infers information from order flow not only within each market but also across markets, as illustrated in Section 2. We now formalize this intuition in the general model.
@@ 4.1 The Posterior and Pricing Kernel
$\widetilde{W}(x_{i}|s_{k})+\varepsilon_{i},\;i=1,2,\cdots.$ (8)
The market maker applies Bayes’ Rule according to his belief about how the insider trades. His unknown parameter is $s_{k}$, with prior probability $\pi_{0}(s_{k})$, $1\leq k\leq K$. His observed data is order flow $\omega$.
Let $\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}$ and $\mathbb{P}_{\!\scriptscriptstyle 0}$ denote the probability densities specifying the stochastic sequences $(\widetilde{W}(x_{i}|s_{k})+\varepsilon_{i})$ and $(\varepsilon_{i})$, respectively, on $\Omega$. In other words, $\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)$ is the likelihood of $\omega$ when the insider’s order is $\widetilde{W}(\,\cdot\,|s_{k})$, while $\mathbb{P}_{\!\scriptscriptstyle 0}(\omega)$ is the likelihood when the insider’s order is zero across markets.
To calculate the conditional likelihood $\mathbb{P}(\omega|s_{k})$ of $\omega$, the market maker takes the likelihood ratio of $\mathbb{P}{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)$ over $\mathbb{P}{\scriptscriptstyle 0}(\omega)$,
$\frac{\mathbb{P}_{\scriptscriptstyle\widetilde{W}(\,\cdot\,|s_{k})}(\omega)}{\mathbb{P}_{\!\scriptscriptstyle 0}(\omega)}=\exp\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right).$ (9)
(For finitely many states, this likelihood ratio simplifies to a ratio of normal densities on a finite-dimensional Euclidean space—e.g., see Equation (1) for the three-state case.)
The market maker then applies Bayes’ Rule to obtain his posterior distribution over $s_{k}\in S$, after receiving order flow $\omega$,
[…]
@@
which is determined by the sufficient statistic
$\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}\right)_{s_{k}\in S}\in\mathbb{R}^{K}.$ (12)
4.2 Sufficient Statistics
Recall the following intuition from Section 2: if the market maker expects the informed demand to be a straddle under high volatility, then receiving order flow consistent with a straddle would lead him to conclude that the true volatility is likely high. His sufficient statistic $\Delta(\omega)$, defined in Equation (2), measures the alignment of order flow with expected insider demands: it is the projection coefficient of order flow $\omega=(\omega_{1},\omega_{2},\omega_{3})$ onto the difference between expected insider demands under high and low volatility, (1,0,1)(0,1,0)(1,0,1)-(0,-1,0). A high $\Delta(\omega)$ indicates that the order flow aligns closely with the straddle $(1,0,1)$, leading the market maker to revise his belief towards high volatility.
In Theorem 4.1, we extend the above intuition to arbitrary information structures. For each signal $s_{k}$, the market maker projects the order flow $\omega$ onto the demand he expects from the insider conditional on $s_{k}$, $\widetilde{W}(\,\cdot\,|s_{k})$. His posterior probability of signal $s_{k}$ is determined by the resulting noise-adjusted projection coefficient:
$\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}}.$ (13)
A high coefficient indicates that order flow $\omega$ aligns closely with the market maker’s belief about informed demand conditional on $s_{k}$, leading him to revise the probability of $s_{k}$ upwards, and vice versa. Thus, the collection (12) of projection coefficients across signals serves as his sufficient statistic.
The market maker adjusts for the noise trading intensities in making cross-market inferences. Higher noise levels reduce the informativeness of his sufficient statistic (12). In the limit where $\sigma_{i}\!\rightarrow\!\infty$ for all $i$, this statistic reduces to the trivial zero statistic, and his posterior (11) reduces to the prior $\pi_{0}$, meaning he infers no information from order flow.
When $X$ is a singleton, Theorem 4.1 simplifies to the market maker’s inference in the familiar single-asset case: if the market maker expects the insider to buy when the asset value is high, receiving a buy order would lead the former to revise the asset price upwards, and vice versa.
For further analysis, it is convenient to encode the coefficient (13) in a definition along with its insider counterpart:

### v39

@@ 1 Introduction
Despite this robust evidence of joint price formation across option contracts, existing informed-trading models (e.g., Back (1993); Biais and Hillion (1994); Brennan and Cao (1996); Easley et al. (1998); Collin-Dufresne et al. (2021)) remain confined to single-option settings and therefore cannot speak to the cross-contract mechanisms at the core of these empirical findings—cross-strike dynamics, implied-volatility-surface formation, and multi-option trades such as the straddle. Our model yields cross-option equilibrium restrictions that sharpen and generalize the empirical intuition while generating additional testable implications. For instance, we identify cross-strike volume imbalance as the equilibrium channel behind the return predictability of signed option volume documented by Pan and Poteshman (2006), and provide the first equilibrium rationale for the volatility-forecasting power of straddle demand observed in Ni et al. (2008) and Goyal and Saretto (2009).
@@
Empirical work consistently finds that access to derivatives markets reduces information asymmetry and improves price efficiency—evidence spans futures (e.g., Garbade and Silber (1983); Chan (1992)), credit derivatives (e.g., Blanco et al. (2005); Longstaff et al. (2005); Acharya and Johnson (2007)), as well as options (e.g., Hu (2018)). Theoretical foundations for this efficiency channel, however, remain limited. Our model isolates a simple and intuitive mechanism: once derivatives are tradable, an informed trader is no longer restricted to generally suboptimal linear demand and can instead express his information through targeted state-contingent positions. In other words, we show that the linear equilibrium in Kyle (1985) is not robust to the introduction of derivatives. [^4] See Example 7.1. Derivatives thus expand the information that prices can aggregate beyond what the underlying asset alone can support.
@@
The reduction of Section 6 is the key modeling innovation that delivers tractability at our full Arrow-Debreu generality and yields an equilibrium characterized by a single endogenous constant. Readers who prefer the high-level overview without the formal developments can see the reduction distilled in Section 6.3, which links directly back to the simplified environment of Section 2.
@@ 2 Basic Intuition
In choosing the scale $\alpha$ of his long-short portfolio, the insider trades off the marginal payoff from exploiting the true payoff distribution against the marginal cost of trading. That cost has two components. First, there is the cost of executing at prevailing prices, absent any price impact. Second, there is an informational component: larger trades make order flow more informative for the market maker, which amplifies cross-market price impact and erodes the insider’s informational advantage. The optimal scale $\alpha^{*}$ balances these margins (see Figure 1a). Equilibrium is characterized by this single endogenous constant—nontrivially, this same equilibrium structure carries over to the general model (see Section 6.3). Allowing correlated noise trades, varying noise intensity, and non-uniform priors has no material effect on the main results. [^7] Correlated Noise Trades Allowing noise trades to be correlated (holding their marginal variances fixed) changes the pattern of uninformed order flow, but not equilibrium price impact at the margin. The reason is that the market maker conditions on the noise covariance when extracting information, so predictable co-movement in noise is effectively filtered out. For example, if noise co-moves in the opposite direction of payoff co-movement, the insider optimally trades more aggressively because order flow is better camouflaged; anticipating this, the market maker adjusts his supply schedules, leaving the marginal price response (the impact matrix) unchanged. This is analogous to the single-asset case where biased noise shifts levels but not the slope of price impact.  [^8] Varying Noise Intensity Suppose noise intensity differs across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$ for $i=1,2,3$. Then the market maker “precision-weights” order flow, and the sufficient statistic becomes $\Delta(\omega)=\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}.$ This is intuitive. A larger $\sigma_{i}$ means order flow in market $i$ is noisier (lower signal-to-noise), so a given trade there is less informative and receives less weight in inference. Consequently, higher noise trading intensity attenuates price impact both within a market and across markets.  [^9] General Prior With a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior odds are just the likelihood ratio from (1) reweighted by the prior odds $\frac{\pi_{1}(s_{1}\mid\omega)}{\pi_{1}(s_{2}\mid\omega)}=\exp\!\big(2\alpha\,\Delta(\omega)\big)\cdot\frac{\pi_{0}(s_{1})}{\pi_{0}(s_{2})},\qquad\Delta(\omega)=\omega_{1}+\omega_{3}-\omega_{2}.$ Intuitively, a higher $\pi_{0}(s_{1})$ means the market maker needs less order-imbalance evidence to attribute trading to the $s_{1}$ “story.” In the toy volatility interpretation, co-buying in the two straddle legs is therefore taken as stronger evidence of high volatility, which amplifies their put-call cross price impact.
@@
If fewer than all three AD securities are available for trade, both the insider’s profit and the information efficiency of prices decline. Here, information efficiency is naturally measured by the posterior probability weight that the market maker assigns to the true signal. Derivative markets allow informed traders’ orders to better express (and therefore reveal) their information, thereby facilitating its incorporation into prices. Therefore, derivatives do more than “add leverage”: they change what information can be traded and how that information propagates across markets. This result provides an equilibrium foundation for a long-held intuition that derivatives concentrate informed trading in directions that are difficult to trade in the underlying alone, and it is consistent with empirical evidence that option markets contain incremental information about underlying returns and higher moments.
Example: Straddle
A key implication is not merely that derivative trades reflect information, but that information is transmitted through cross-market price impact. We demonstrate this mechanism in the case of volatility. Interpret the three states $x_{1}<x_{2}<x_{3}$ as the possible $t=1$ realizations of an underlying asset price, and normalize the $t=0$ price to be $x_{2}$. Then $s_{1}$ is the “high volatility” signal and $s_{2}$ the “low volatility” signal. The signal $s_{1}$ reveals that the asset price will move, but without a specified direction, whereas $s_{2}$ reveals that there will be no price movement. Security $x_{1}$ is a put-like claim (it pays when the asset price falls), while $x_{3}$ is call-like (it pays when the asset price rises). If the insider observes $s_{1}$, he buys the put-call pair $(x_{1},x_{3})$—a “straddle”—to capitalize price movement regardless of direction. In real option markets, the straddle is a standard volatility strategy (see Hull (2015)) and is widely discussed in the empirical literature. [^10] See Coval and Shumway (2001); Pan and Poteshman (2006); Ni et al. (2008); Driessen et al. (2009); Goyal and Saretto (2009). This strategy, along with other common option strategies, will be derived as an equilibrium portfolio in the general model.
@@ 3 General Model
We place no restrictions on $\eta(\,\cdot\,|\,\cdot\,)$, i.e., no restrictions on the probability distributions across states or the nature of private information.
@@
The general economic intuition we bring out is robust to whether the state and signal spaces are discrete or continuous. A continuous setup—such as the normal prior/payoff environment of Kyle (1985)—can be approximated arbitrarily well by a discrete signal-state grid. We consider the discrete model here to focus on the economic results and minimize technicality. A fully rigorous continuous-space treatment (allowing general priors and signal structures) is provided in Keller and Tseng (2026).
In this single-asset special case, the price impact is the slope $\lambda$ of the market maker’s inverse supply curve $P(Q)=v_{0}+\lambda Q$, as specified by Equations (5) and (6). Price impact is proportional to the (noise-adjusted) signal variation $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$. Higher signal variation leads to greater informed demand variation across signals, making order flow more informative and resulting in a higher price impact. In other words, price impact is determined by the insider’s information intensity. As suggested in Section 2, this intuition generalizes across AD markets.
@@
The AD formulation can be readily specialized to option markets. Suppose the states $x_{1}<x_{2}<\cdots$ are the possible $t=1$ prices of an underlying asset. Let $x_{i_{0}}$ denote the market maker’s prior mean of the $t=1$ price at $t=0$. Then for any state-contingent portfolio $W(\,\cdot\,)$ there exist holdings $(a_{i})$ such that
$W(x_{j})$ $=a_{i_{0}}\,(x_{j}-x_{i_{0}})+\sum_{i<i_{0}}a_{i}\,(x_{j}-x_{i})_{-}+\sum_{i>i_{0}}a_{i}\,(x_{j}-x_{i})_{+},\qquad j=1,2,\ldots.$ (7)
This is the standard (discrete) Breeden–Litzenberger formula (see Back (2010), Exercise 3.5). Thus, $W(\,\cdot\,)$ can be implemented by holding $a_{i_{0}}$ shares of the underlying, $a_{i}$ puts at strikes $x_{i}<x_{i_{0}}$, and $a_{i}$ calls at strikes $x_{i}>x_{i_{0}}$. So the AD securities can be replaced by the underlying plus a menu of options.
In this options representation, the insider observes $s_{k}$ and hence the signal-implied payoff distribution $\eta(\,\cdot\,\mid s_{k})$ for the underlying, and submits orders $(a_{i})$ across the underlying and options. In each market $i$, aggregate order flow is
$\omega_{i}=a_{i}+\varepsilon_{i},\quad\mbox{ with noise trades }\varepsilon_{i}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\sigma_{i}^{2}).$
The market maker prices contracts conditional on $\omega=(\omega_{i})$ exactly as in the AD formulation. Accordingly, all results derived in the AD formulation translate directly to options.
@@
The options formulation relates to the single-asset case simply as follows:
@@
(i) If signals shift the payoff distribution only through its mean, and trading is restricted to the underlying (no options), then our model collapses to the Kyle (1985) setting.
(ii) If all signal-implied payoff distributions share the same mean, then the insider does not trade the underlying at all. In other words, exploiting information beyond the expected payoff requires options.
Proposition 3.3(i) is a knife-edge case—once options are available, the linear equilibrium (5) in Kyle (1985) is not robust. Options let the insider concentrate on the strikes where his signal matters and avoid low-probability tail exposure, so equilibrium informed demand is inherently nonlinear. We will illustrate this explicitly in Example 7.1.
@@ 4 Cross-Market Inference
The market maker’s inference problem is infinite-dimensional: as noisy order flow arrives across many claims, he must infer the informed portfolio behind it, which is unrestricted. Making the underlying economic intuition precise therefore requires going beyond finite-dimensional Bayesian updating. Once we do, the takeaway is simple. His sufficient statistic is the degree to which order flow aligns with the insider-demand portfolio he expects under each signal—a direct generalization of the cross-market imbalance $\Delta(\omega)$ of Section 2.
@@ 4.1 The Posterior and Pricing Kernel
Formally, order flow is a (random) sequence $\omega=(\omega_{i})_{i=1,2,\ldots}$ taking values in the space $\Omega$ of countable sequences. Under the market maker’s belief $W(\,\cdot\,|\,\cdot\,)$ about how the insider trades, conditional on signal $s_{k}$ the observed order flow is
$\omega_{i}\,=\,W(x_{i}|s_{k})+\varepsilon_{i},\qquad i=1,2,\ldots,$ (8)
where $\varepsilon_{i}\stackrel{{\scriptstyle d}}{{\sim}}\mathcal{N}(0,\sigma_{i}^{2})$ are noise trades.
The market maker treats $s_{k}$ as the latent parameter, with prior $\pi_{0}(s_{k})$ for $k=1,\ldots,K$, and updates from the realized order flow $\omega$. Let $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ denote the likelihood of $\omega$ under (8) when the insider’s mean order is $W(\,\cdot\,|s_{k})$, and let $\mathbb{P}_{\!\scriptscriptstyle 0}$ denote the noise-only likelihood (i.e., $W\equiv 0$). Since the noise-only likelihood does not depend on signal, the market maker can summarize the likelihood of $\omega$ conditional on $s_{k}$ via the likelihood ratio [^13] Formally, $\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}$ and $\mathbb{P}_{\!\scriptscriptstyle 0}$ are probability measures on $\Omega$. The likelihood ratio (9) is their Radon-Nikodym derivative.
$\frac{\mathbb{P}_{\scriptscriptstyle W(\,\cdot\,|s_{k})}(\omega)}{\mathbb{P}_{\!\scriptscriptstyle 0}(\omega)}\,=\,\exp\!\left(\sum_{i}\frac{W(x_{i}|s_{k})\,\omega_{i}}{\sigma_{i}^{2}}-\frac{1}{2}\sum_{i}\frac{W(x_{i}|s_{k})^{2}}{\sigma_{i}^{2}}\right).$ (9)
By Bayes’ Rule, after observing order flow $\omega$ the market maker’s posterior belief over $s_{k}\in S$ is then
@@
with the proportionality constant $C(\omega)$ normalizing the expression to sum to one over $S$. Given this posterior, the market maker prices AD claims by the zero-profit condition (3).
@@
[…]
$\left(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\,\omega_{i}}{\sigma_{i}^{2}}\right)_{s_{k}\in S}\in\mathbb{R}^{K},$ (11)
that is, how closely the realized order flow aligns with the expected signal-contingent demand profiles. These scores therefore form the sufficient statistic for his inference.
4.2 Sufficient Statistic
Recall the three-state illustration in Section 2. There, the market maker does not treat $(\omega_{1},\omega_{2},\omega_{3})$ as three standalone order imbalances. What matters for inference is not whether one particular contract has buy pressure, but whether the order flow across contracts resembles a recognizable informed portfolio. In that toy volatility interpretation, he effectively asks: does the joint order flow look more like the portfolio he expects under high volatility (a straddle) or under low volatility? The cross-market imbalance $\Delta(\omega)$ of (2) summarizes exactly this “fit.”
Theorem 4.1 makes this intuition precise under arbitrary information structures. For each signal $s_{k}$, the market maker computes a precision-weighted alignment score
$\sum_{i}\frac{\widetilde{W}(x_{i}|s_{k})\,\omega_{i}}{\sigma_{i}^{2}},$ (12)
which measures how well the realized order flow matches the demand profile he expects conditional on $s_{k}$. A high score means: “the order flow looks like the portfolio I would expect if traders were acting on $s_{k}$.”
His posterior is therefore determined by the vector of alignment scores (11)—one for each information type. The weights $1/\sigma_{i}^{2}$ are the signal-to-noise adjustment: trades in noisier markets (larger $\sigma_{i}$) are less informative and therefore receive less weight in inference. At the extreme, if $\sigma_{i}\to\infty$ for all $i$, order flow becomes pure noise, the alignment scores shrink to zero, and the posterior reduces to the prior.
When $X$ is a singleton, Theorem 4.1 reduces to the familiar Kyle single-asset inference reviewed in Section 3: the sufficient statistic (11) becomes a single scalar, a scaled version of net signed order flow $\omega$. Accordingly, buy pressure makes high values more likely and raises the market maker’s posterior mean (the price), while sell pressure does the opposite.
We will refer to (12) as the market maker’s alignment score for signal $s_{k}$. We will define the analogous impact exposure when $\omega$ is replaced by a candidate insider portfolio $W$, since that quantity will govern the portfolio’s price impact in Section 5.